OpenStudy (anonymous):

An ice cream cone has a diameter of 2 inches and a height of 5 inches. What is the volume of the ice cream cone? Round your answer to the nearest hundredth. (Use 3.14 for π.)

OpenStudy (anonymous):

\[V_{Cone} = \frac{1}{3}πr^{2}h\]\[r = \frac d2\]

OpenStudy (anonymous):

So, using those two formulas: r = d/2 r = 2/2 r = 1 Now use the volume formula. \[V_{Cone} = \frac 13π1^{2}(5)\]\[V_{Cone} = \frac{5π}3\]Can you simplify this?

OpenStudy (anonymous):

5.23 ?

OpenStudy (anonymous):

Pretty much. 5.235 turns to 5.24 since you're rounding to the tenth place.

OpenStudy (anonymous):

thank youuu

OpenStudy (anonymous):

np :)